OpenStudy (anonymous):

the equation of line CD is y=3x-3. write an equation ofa line perpendicular to line CD in slope-intercept form that conatins point (3,1). a) y=3x+0 b) y=-3x-8 c) y=-1/3x+2 d) y=1/3x +0 medal and fan will be given

OpenStudy (anonymous):

@ganeshie8 hi can you help

OpenStudy (anonymous):

@paki hey can you help me

OpenStudy (anonymous):

@zepdrix hi can you help me please

OpenStudy (anonymous):

ok so your like is currently in slope-intercept form (y=mx+b) m=slope so the slope of your original line would be what?

OpenStudy (anonymous):

isnt the slope 3

OpenStudy (anonymous):

yes perfect

OpenStudy (anonymous):

so 3 can also be written as positive \[\dfrac{3}{1}\]

OpenStudy (anonymous):

to find the slope of a perpendicular line, all you need to do is flip those and change the sign so it becomes negative \[\dfrac{1}{3}\]

OpenStudy (anonymous):

so would it be -1/3x+2

OpenStudy (anonymous):

from there, you can put in your two points into point slope form \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

yes it would :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

sure thing :)